OpenStudy (anonymous):

A 70g piece of ice at 0 Celsius is added to a sample of water at 6 Celsius. All of the ice melts and the temp of the water decreases to 0 Celsius. How many grams of water were in the sample.

OpenStudy (anonymous):

specific heat equation mC(dt)=mC(dt) dt = t_f-t_o m is mass C is specific heat

OpenStudy (anonymous):

http://hyperphysics.phy-astr.gsu.edu/hbase/thermo/spht.html Q must be equal to each other

OpenStudy (anonymous):

I'm still confused. Surprise, surprise! This is like a foreign language to me. I started out by multiplying 70g by 80.cal/1g.ice=5600 cal. Is this even close? I know it's not the only step.

OpenStudy (anonymous):

C does not mean call C is the specific heat of the material you need to look this up the specific heat of ice and water are different

OpenStudy (anonymous):

specific heat of water C_water= 4.186 joule/gram °C specific heat of ice C_ice =2.108 kJ/kg-K \[m_{ice}C_{ice}\Delta t=m_{water}C_{water}\Delta t\]

OpenStudy (anonymous):

ice goes from 0 to 6 so delta t = 6-0 water goes from 6 to 0 so delta t = 0-6 sorry forgot a negative sign in there

OpenStudy (anonymous):

Ok, so is this close? 70g*2.108kj*6=m*1.0*-6

OpenStudy (anonymous):

70*2.108*6=m*4.186*6 now solve for m

OpenStudy (anonymous):

35.25g??

OpenStudy (anonymous):

Well, that answer is wrong, so obviously I have NO idea what on earth i'm doing.

OpenStudy (anonymous):

Or you forgot the significant digit! It may 35,3g

OpenStudy (anonymous):

Nope, it's still wrong. :(